IF A+B =23 AND AB=102 FIND A2+B2   AND ALSO FIND ALL THE REAL VALUES OF A AND B

Grade:9

1 Answers

Aryan Maurya
6 years ago
a +b=23(A+b) ^2=a*a +b*b +2abA23*23=a*a+b*b+204529-204=a^2 +b^2=325( A-b)^2=a*a +b*b - 2ab(A-b)^2=325-204=121Square rooting both sides A-b=11 (i)A+b=23 (ii)Equation i +ii2a=34a=17B=6